### v1

@@ 4.3 Ablation Study
In this section, we further analyze the interplay between our proposed optimization strategy and perception performance to address two questions: 1) How does our optimization strategy inform LiDAR placement to improve robustness against various forms of corruption? 2) How can our optimization strategy enhance perception performance in scenarios with limited placement options?
Optimizing Performance under Corruption. In the previous subsection, we compared the performance of the optimized placement with existing baseline placements under both clean and adverse conditions (cf. Figs. 5(c), 5(d), 5(a) and 5(b)). Despite outperforming the baseline design under corruption, the optimized placement still fell short of the clean setting’s performance. To address this, we conduct placement optimization directly on the adverse data distribution for LiDAR semantic segmentation. Specifically, we perform our M-SOG-based optimization approach utilizing the P-SOG derived from corrupted conditions.
Quantitatively, the configurations derived from optimization on adverse data outperform the configurations generated solely on clean data for different baseline methods. Specifically, for SPVCNN, the configuration optimized on adverse data achieves a significantly higher performance under corruption, with a mIoU of 51.1, compared to the configurations optimized solely on clean data (48.6 mIoU). Similarly, for MinkUNet and Cylinder3D, we observe slight performance gains when evaluating corrupted settings by optimizing placement on corrupted data (from 48.6 mIoU to 49.8 mIoU for MinkUNet; from 47.3 mIoU to 47.9 mIoU for Cylinder3D). These results indicate that customizable optimization tailored for adverse settings is an effective approach to enable robust perception.
(a) Clean (Plane)
(b) Corruptions (Plane)
Figure 6: Optimization results with constrained placement options. From left to right are Trapezoid, Line, Square, and our optimized placement on 2D plane.
Optimizing under Constrained Placement Options. In automotive design practice, placing LiDARs at different heights on the roof might affect aesthetics and aerodynamics. Therefore, we investigate the effect of our optimization algorithm in the presence of constraints, by limiting the LiDARs to the same height and considering only 2D placements. We fix the height of all LiDARs to find the optimal placement on the horizontal plane of the vehicle roof. We use the line, square, and trapezoid placements in Fig. 4 for comparison. Experimental results in Figs. 6(a) and 6(b) show that our optimized placement generally outperforms baseline placement at the same LiDAR height on both clean and corruption settings.
@@ 5 Conclusion
In this work, we presented Place3D, a comprehensive and systematic LiDAR placement evaluation and optimization framework. We introduced the Surrogate Metric of Semantic Occupancy Grids (M-SOG) as a novel metric to assess the impact of various LiDAR placements. Building on this metric, we culminated an optimization approach for LiDAR configuration that significantly enhances detection and segmentation performance. We validate the effectiveness of our optimization approach through a multi-condition multi-LiDAR point cloud dataset and establish a comprehensive benchmark for evaluating both baseline and our optimized LiDAR placements on detection and segmentation in diverse conditions. Our optimized placements demonstrate superior robustness and perception capabilities, outperforming all baseline configurations. By shedding light on refining the robustness of LiDAR placements for both detection and segmentation tasks under adverse conditions, we anticipate that our work will pave the way for further advancements in this field.
Acknowledgment
This work was supported by the Office of Naval Research (Grant #: N00014-24-1-2137; Program Manager: Michael “Q” Qin).
@@ 6 Place3D Dataset
6.1 Statistics
Our dataset consists of a total of eleven LiDAR placements, in which seven baselines are inspired by existing self-driving configurations from autonomous vehicle companies and four LiDAR placements are obtained by optimization. Each LiDAR placement contains four LiDAR sensors. For each LiDAR configuration, the sub-dataset consists of 13,600 frames of samples, comprising 11,200 samples for training and 2,400 samples for validation, following the split ratio used in nuScenes [13]. We combined every 40 frames of samples into one scene, with a time interval of 0.5 seconds between each frame sample.
Table 6: Comparisons between Place3D and popular driving perception datasets with supports for the LiDAR semantic segmentation and 3D object detection tasks. Notably, the classes reported below represent those used for training and testing, with the actual number potentially exceeding what is shown.
@@ 6.1 Statistics
6.2 Data Collection
6.2.1 Simulator Setup.
We choose four maps (Towns 1, 3, 4, and 6, c.f. Fig. 7) in CARLA v0.9.10 to collect point cloud data and generate ground truth information. For each map, we manually set 6 ego-vehicle routes to cover all roads with no roads overlapped. The frequency of the simulation is set to 20 Hz.
6.2.2 LiDAR Setup.
LiDAR point cloud data is collected every 0.5 simulator seconds utilizing the Semantic LIDAR sensor in CARLA. The attributes of Semantic LIDAR sensor are listed in Tab. 7. Notably, we utilize LiDARs with relatively low resolution to increase the challenge in object detection and semantic segmentation, leading to lower scores than those on the nuScenes [13] leaderboard. This is based on the observation that LiDAR point clouds become sparse when detecting distant objects and the use of multiple LiDARs is specifically to enhance the perception of sparse point clouds at long distances in practical applications. Our experiments with low-resolution LiDARs allow for a more evident assessment of the impact of LiDAR placements on the perception of sparse point clouds.
(a) Town 1
(b) Town 3
(c) Town 4
(d) Town 6
Figure 7: The maps used to collect Place3D data in CARLA v0.9.10.
(a) fog
(b) wet ground
(c) snow
(d) motion blur
(e) crosstalk
(f) incomplete echo
(a) fog
(b) wet ground
(c) snow
(d) motion blur
(e) crosstalk
(f) incomplete echo
Table 8: Summary of the semantic categories defined in the Place3D dataset.
@@ 6.2 Data Collection
Pedestrian $\blacksquare$ 4 Humans that walk, ride, or drive any kind of vehicle or mobility system, such as bicycles, scooters, skateboards, wheelchairs, etc.
@@
6.3 Label Mappings
6.3.1 LiDAR Semantic Segmentation.
There are a total of 21 semantic labels in our LiDAR semantic segmentation dataset, which are 1Building, 2Fence, 3Other, 4Pedestrian, 5Pole, 6Road Line, 7Road, 8Sidewalk, 9Vegetation, 10Vehicle, 11Wall, 12Traffic Sign, 13Ground, 14Bridge, 15Rail Track, 16Guard Rail, 17Traffic Light, 18Static, 19Dynamic, 20Water, and 21Terrain. Tab. 8 presents the detailed definition of each semantic class in the Place3D dataset. Additionally, we include a Unlabeled tag to denote elements that have not been categorized.
6.3.2 3D Object Detection.
@@ 6.3 Label Mappings
6.4 Adverse Conditions
As shown in Fig. 8 and Fig. 9, the adverse conditions in the Place3D dataset can be categorized into three distinct groups: 1) Severe weather conditions, including “fog”, which causes back-scattering and attenuation of LiDAR points due to water particles in the air; “snow”, where snow particles intersect with laser beams, affecting the beam’s reflection and causing potential occlusions; and “wet ground”, where laser pulses lose energy upon hitting wet surfaces, leading to attenuated echoes. 2) External disturbances like “motion blur”, which resulted from vehicle movement that causes blurring in the data, especially on bumpy surfaces or during rapid turns. 3) Internal sensor failure, such as “crosstalk”, where the light impulses from multiple sensors interfere with each other, creating noisy points; as well as “incomplete echo”, where dark-colored objects result in incomplete LiDAR readings.
6.5 License
@@
7 Additional Implementation Detail
In this section, we present additional implementation details to facilitate the reproduction of this work.
7.1 LiDAR Placement
We adopt five commonly employed heuristic LiDAR placements, which have been adopted by several leading autonomous driving companies, as our baseline. We create another two baseline placements by adjusting the rolling angles of LiDAR sensors. We obtain two placements in optimization and two placements in the ablation study. All configurations are presented in Tab. 9.
Table 9: The configuration coordinates of different LiDAR placement strategies in this work with respect to their ego-vehicle coordinate frames from four LiDAR sensors.
@@ 7.3 Hyperparameters
The detailed training configurations of the four LiDAR semantic segmentation models, i.e., MinkUNet [21], SPVCNN [78], PolarNet [99], and Cylinder3D [105], are presented in Tab. 10. The detailed training configurations of the four 3D object detection models, i.e., PointPillars [50], CenterPoint [95], BEVFusion-L [58], and FSTR [97], are presented in Tab. 11.
@@
8 Additional Quantitative Result
In this section, we provide additional quantitative results, such as the class-wise LiDAR semantic segmentation results, to better support the findings and conclusions drawn in the main body of this paper.
@@ 8 Additional Quantitative Result
$\bullet$ Placement: 2D Plane
@@
8.1 LiDAR Semantic Segmentation
In this section, we provide the complete (i.e., the class-wise IoU scores) results for LiDAR semantic segmentation under different LiDAR placement strategies.
We showcase the per-class LiDAR semantic segmentation results of MinkUNet [21], SPVCNN [78], PolarNet [99], and Cylinder3D [105] in Tab. 12. The performance of these methods is evaluated under different LiDAR placement strategies. Different LiDAR placement strategies demonstrated varying propensities towards particular classes. For instance, Pyramid performs well for building, guard rail, and vegetation compared with other placements, possibly due to an increased vertical field of view that captures these taller or layered structures more effectively. Ours provides the generally best balance of performance across categories, with high scores in building, road, and vehicle suggesting an effective all-around coverage for various object types. The complete benchmark results of LiDAR semantic segmentation under clean and adverse conditions are presented in Tab. 13.
We showcase the complete results of the ablation study for semantic segmentation in Tab. 15. we first explore the performance of our optimization algorithm under placement constraints, specifically, standardize the elevation of LiDARs to analyze their optimal positioning on the vehicle’s roof’s horizontal plane. Our optimized placement 2D plane achieved better performance compared with line, square, and trapezoid configurations. Further, we optimize the placement for adverse conditions. While the observation suggests a trade-off in clean data compared with the Ours, the configuration optimized on corruption data (Corruption optimized) achieves a significantly higher performance in adverse conditions than both baselines in Tab. 13 and Ours, indicating that custom optimization strategies, tailored for challenging conditions, represent an effective methodology to enhance robust perception capabilities in adverse environments.
8.2 3D Object Detection
In this section, we present more experimental results for the 3D robustness evaluation of 3D object detection under adverse conditions. The performance of PointPillars[50], CenterPoint[95], BEVFusion-L[58], and FSTR[97] is evaluated under different LiDAR placement strategies in Tab. 14. Across all placements and conditions, there is a trend where all methods suffer a drop in performance in adverse conditions compared to clean conditions, highlighting the challenge that adverse conditions pose to 3D object detection. While some placements, like Pyramid and Square, appear to maintain relatively high performance under both clean and adverse conditions, the Ours placement has the highest average mAP under adverse conditions. In addition, certain placements exhibit relative strengths against specific types of corruption, which can inform the development of more resilient object detection systems for autonomous vehicles. For instance, Pyramid and Pyramid-Roll, appear to handle fog better than others, possibly due to a configuration that captures a more diverse set of angles which could mitigate the scattering effect of fog on LiDAR beams.
9 Additional Qualitative Result
9.1 Clean Condition
We showcase the qualitative results of the MinkUNet [21] model using our optimized LiDAR placement. As shown in Fig. 10, the depiction across the first four rows demonstrates a commendable performance of our LiDAR placement strategy during clean conditions. This is evidenced by the predominance of gray in the error maps, which indicate a high rate of correct predictions. Also, the fifth row presents a notable exception, illustrating a failure case. This particular scene exhibits a heightened complexity with a diverse array of objects and potential occlusions that challenge predictive accuracy. The qualitative results underscore the importance of continuous refinement of LiDAR placement to achieve better perception across a comprehensive spectrum of scenarios.
9.2 Adverse Conditions
In Fig. 11, we showcase the qualitative results of MinkUNet [21] under different adverse conditions utilizing our optimized LiDAR placement. The corruptions from top to bottom rows are “fog”, “wet ground”, “motion blur”, “crosstalk”, and “incomplete echo”. As can be seen, the LiDAR semantic segmentation model encounters extra difficulties when predicting under such scenarios. Erroneous predictions tend to appear in regions contaminated by different types of noises, such as airborne particles, disturbed laser reflections, and jittering scatters. It becomes apparent that enhancing the model’s robustness under these adverse conditions is crucial for the practical usage of driving perception systems. To achieve better robustness, we design suitable LiDAR placements that can mitigate the degradation caused by corruptions. As discussed in Tab. 13, Tab. 14, and Tab. 15, our placements can largely enhance the robustness of various models.
Figure 10: Qualitative assessments of the MinkUNet [21] model using our LiDAR placement strategy. The model is tested under the clean condition. The error maps show the correct and incorrect predictions in gray and red, respectively. Kindly refer to Tab. 8 for color maps. Best viewed in colors and zoomed-in for details.
Figure 11: Qualitative assessments of the MinkUNet [21] model using our LiDAR placement strategy. The model is tested under adverse conditions. The error maps show the correct and incorrect predictions in gray and red, respectively. Kindly refer to Tab. 8 for color maps. Best viewed in colors and zoomed-in for details.
10 Proofs

### v2

@@ 4.3 Ablation Study
In this section, we further analyze the interplay between our proposed optimization strategy and perception performance to address two questions: 1) How does our optimization strategy inform LiDAR placements to improve robustness against various forms of corruption? 2) How can our optimization strategy enhance perception performance in scenarios with limited placement options?
Optimizing Against Corruptions. In the previous subsection, we deploy the placement optimized on the clean point clouds directly on corrupted ones to assess the robustness. To further showcase the capability of our method against corruptions, we compute the M-SOG scores utilizing the P-SOG derived from the corrupted semantic point clouds under adverse conditions and perform the optimization process, as shown in Figure 6(a). Quantitatively compared with Ours in Table 2, the configurations derived from optimization on adverse data outperform the configurations generated solely on clean data. These results indicate that customizable optimization tailored for adverse settings is an effective approach to enable robust 3D perception.
Optimizing 2D Placements. Placing LiDARs at different heights on the roof might affect automotive aesthetics and aerodynamics. Therefore, we investigate the effect of our optimization algorithm in the presence of constraints, by limiting the LiDARs to the same height and considering only 2D placements. We fix the height of all LiDARs to find the optimal placement on the horizontal plane of the vehicle roof. We use the Line, Square, and Trapezoid placements in Figure 2 for comparison. Experimental results in Figures 6(b) and 6(c) show that our optimized LiDAR placements outperform baseline placements at the same LiDAR height on both clean and corruption settings.
Influence of LiDAR Roll Angles. Fine-tuning the orientation angles of the LiDARs on autonomous vehicles is an effective strategy to minimize blind spots and broaden the perception range. For 3D segmentation, Pyramid-roll slightly outperforms Pyramid, whereas Line-roll falls short of Line, which is highly aligned with scores of M-SOG in Table 1. For 3D detection tasks, the situation is reversed, but still consistent with the M-SOG results. This suggests that adjusting the LiDAR’s angle can have varying impacts on the performance of 3D object detection and LiDAR semantic segmentation.
@@ 5 Conclusion
In this work, we presented Place3D, a comprehensive and systematic LiDAR placement evaluation and optimization framework. We introduced the Surrogate Metric of Semantic Occupancy Grids (M-SOG) as a novel metric to assess the impact of various LiDAR placements. Building on this metric, we culminated an optimization approach for LiDAR configuration that significantly enhances LiDAR semantic segmentation and 3D object detection performance. We validate the effectiveness of our optimization approach through a multi-condition multi-LiDAR point cloud dataset and establish a comprehensive benchmark for evaluating both baseline and our optimized LiDAR placements on detection and segmentation in diverse conditions. Our optimized placements demonstrate superior robustness and perception capabilities, outperforming all baseline configurations. By shedding light on refining the robustness of LiDAR placements for both tasks under diverse driving conditions, we anticipate that our work will pave the way for further advancements in this field of research.
@@ Appendix A A The Place3D Dataset
A.1 Statistics
Our dataset consists of a total of eleven LiDAR placements, in which seven baselines are inspired by existing self-driving configurations from autonomous vehicle companies, and four LiDAR placements are obtained by optimization. Each LiDAR placement contains four LiDAR sensors. For each LiDAR configuration, the sub-dataset consists of 13,600 frames of samples, comprising 11,200 samples for training and 2,400 samples for validation, following the split ratio used in nuScenes [7]. We combined every 40 frames of samples into one scene, with a time interval of 0.5 seconds between each frame sample.
A.2 Data Collection
In this section, we elaborate on more details of the simulator setup and LiDAR setup procedures used in collecting the Place3D dataset.
(a) Town 1
(b) Town 3
(c) Town 4
(d) Town 6
Figure 7: The maps used to collect the Place3D data in CARLA v0.9.10.
Table 4: Comparisons between the established benchmark and popular driving perception datasets with support for both the LiDAR semantic segmentation and 3D object detection tasks.
@@ A.2 Data Collection
(a) Waymo
(b) Motional
(c) Cruise
(d) Pony.ai
(e) Zoox
(f) Toyota
(g) Momenta
(h) Ford
Figure 8: A diverse spectrum of existing multi-LiDAR placements employed by major autonomous vehicle companies [91, 72, 22, 76, 119, 89, 71, 27]. Images adopted from original websites.
A.2.1 Simulator Setup
We choose four maps (Towns 1, 3, 4, and 6, cf. Figure 7) in CARLA v0.9.10 to collect point cloud data and generate ground truth information. For each map, we manually set six ego-vehicle routes to cover all roads with no roads overlapped. The frequency of the simulation is set to 20 Hz.
A.2.2 LiDAR Setup
LiDAR point cloud data is collected every 0.5 simulator seconds utilizing the Semantic LIDAR sensor in CARLA. The attributes of Semantic LIDAR sensor are listed in Table 5. Notably, we utilize LiDARs with relatively low resolution to increase the challenge in object detection and semantic segmentation, leading to lower scores than those on the nuScenes [7] leaderboard. This is based on the observation that LiDAR point clouds become sparse when detecting distant objects and the use of multiple LiDARs is specifically to enhance the perception of sparse point clouds at long distances in practical applications. Our experiments with low-resolution LiDARs allow for a more evident assessment of the impact of LiDAR placements on the perception of sparse point clouds.
Table 6: Summary of the semantic categories and their semantic coverage in the Place3D dataset.
@@ A.2.2 LiDAR Setup
Pedestrian $\blacksquare$ 4 Humans that walk, ride, or drive any kind of vehicle or mobility system, such as bicycles, scooters, skateboards, wheelchairs, etc..
@@
A.3 Label Mappings
In this section, we provide more details for the class definitions of the LiDAR semantic segmentation and 3D object detection tasks.
A.3.1 LiDAR Semantic Segmentation
There are a total of 21 semantic categories in our LiDAR semantic segmentation dataset, which are 1Building, 2Fence, 3Other, 4Pedestrian, 5Pole, 6Road Line, 7Road, 8Sidewalk, 9Vegetation, 10Vehicle, 11Wall, 12Traffic Sign, 13Ground, 14Bridge, 15Rail Track, 16Guard Rail, 17Traffic Light, 18Static, 19Dynamic, 20Water, and 21Terrain. Table 6 presents the detailed definition of each semantic class in Place3D. Additionally, we include a Unlabeled tag to denote elements that have not been categorized.
A.3.2 3D Object Detection
@@ A.3.1 LiDAR Semantic Segmentation
A.4 Adverse Conditions
As shown in Figure 9 and Figure 10, the adverse conditions in the Place3D dataset can be categorized into three distinct groups:
  • Severe weather conditions, including “fog”, which causes back-scattering and attenuation of LiDAR points due to water particles in the air; “snow”, where snow particles intersect with laser beams, affecting the beam’s reflection and causing potential occlusions; and “wet ground”, where laser pulses lose energy upon hitting wet surfaces, leading to attenuated echoes. These are commonly occurring corruptions in real driving conditions.
  • External disturbances like “motion blur”, which results from vehicle movement that causes blurring in the data, especially on bumpy surfaces or during rapid turns.
  • Internal sensor failure, such as “crosstalk”, where the light impulses from multiple sensors interfere with each other, creating noisy points; as well as “incomplete echo”, where dark-colored objects result in incomplete LiDAR readings.
A.5 License
@@
Appendix B B Additional Implementation Detail
(a) fog
(b) wet ground
(c) snow
(d) motion blur
(e) crosstalk
(f) incomplete echo
(a) fog
(b) wet ground
(c) snow
(d) motion blur
(e) crosstalk
(f) incomplete echo
In this section, we present additional implementation details to facilitate the reproduction of the data generation, placement optimization, and performance evaluations in this work.
B.1 LiDAR Placement
We adopt five commonly employed heuristic LiDAR placements, which have been adopted by several leading autonomous driving companies (see Figure 8), as our baseline. We create another two baseline placements by adjusting the rolling angles of LiDAR sensors. We obtain two placements in optimization and two placements in the ablation study. All configurations are presented in Table 7.
Table 7: The configuration coordinates of different LiDAR placement strategies in this work with respect to their ego-vehicle coordinate frames from the four LiDAR sensors.
@@ B.3 Hyperparameters
The detailed training configurations of the four LiDAR semantic segmentation models, i.e., MinkUNet [18], SPVCNN [86], PolarNet [111], and Cylinder3D [117], are presented in Table 8. The detailed training configurations of the four 3D object detection models, i.e., PointPillars [51], CenterPoint [107], BEVFusion-L [63], and FSTR [109], are presented in Table 9.
@@
Appendix C C Additional Quantitative Result
In this section, we provide additional quantitative results to better support the findings and conclusions drawn in the main body of this paper.
C.1 LiDAR Semantic Segmentation
We provide the complete (i.e., the class-wise IoU scores) results for LiDAR semantic segmentation under different LiDAR placement strategies.
We showcase the per-class LiDAR semantic segmentation results of MinkUNet [18], SPVCNN [86], PolarNet [111], and Cylinder3D [117] in Table 11. The performance of these methods is evaluated under different LiDAR placement strategies. Different LiDAR placement strategies demonstrated varying propensities towards particular classes. For instance, Pyramid performs well for building, guard rail, and vegetation compared with other placements, possibly due to an increased vertical field of view that captures these taller or layered structures more effectively. Ours provides the generally best balance of performance across categories, with high scores in building, road, and vehicle suggesting an effective all-around coverage for various object types. The complete benchmark results of LiDAR semantic segmentation under clean and adverse conditions are presented in Table 12. We showcase the correlation between M-SOG and LiDAR semantic segmentation performance under clean and adverse conditions in Figure 11 and 12.
We showcase the complete results of the ablation study for semantic segmentation in Table 14. we first explore the performance of our optimization algorithm under placement constraints, specifically, standardize the elevation of LiDARs to analyze their optimal positioning on the vehicle’s roof’s horizontal plane. Our optimized placement 2D plane achieved better performance compared with line, square, and trapezoid configurations. Further, we optimize the placement for adverse conditions. While the observation suggests a trade-off in clean data compared with the Ours, the configuration optimized on corruption data (corruption optimized) achieves a significantly higher performance in adverse conditions than both baselines in Table 12 and Ours, indicating that custom optimization strategies, tailored for challenging conditions, represent an effective methodology to enhance robust perception capabilities in adverse environments.
C.2 3D Object Detection
In this section, we present more experimental results for the 3D robustness evaluation of 3D object detection under adverse conditions. The performance of PointPillars [51], CenterPoint [107], BEVFusion-L [63], and FSTR [109] is evaluated under different LiDAR placement strategies in Table 13. Across all placements and conditions, there is a trend where all methods suffer a drop in performance in adverse conditions compared to clean conditions, highlighting the challenge that adverse conditions pose to 3D object detection. While some placements, like Pyramid and Square, appear to maintain relatively high performance under both clean and adverse conditions, the Ours placement has the highest average mAP under adverse conditions. In addition, certain placements exhibit relative strengths against specific types of corruption, which can inform the development of more resilient object detection systems for autonomous vehicles. For instance, Pyramid and Pyramid-Roll, appear to handle fog better than others, possibly due to a configuration that captures a more diverse set of angles which could mitigate the scattering effect of fog on LiDAR beams. We showcase the correlation between M-SOG and 3D Object Detection performance under clean and adverse conditions in Figure 13 and 14.
@@ C.1 LiDAR Semantic Segmentation
C.3 Compare to Existing Approaches
The effectiveness of sensor placement metrics is determined by the linear correlation between these metrics and actual performance. As shown in Table 10, Jiang’s method [40] is applied to roadside placements, while comparable on-vehicle methods, S-MIG [34] and S-MS [57], follow the same process for LiDARs. All these methods use 3D bounding boxes as priors to understand the 3D scene distribution. However, this approach leads to significant deviations from the actual physical boundaries of objects and fails to account for occlusion relationships between objects and the environment in LiDAR applications. Consequently, they cannot accurately describe the scene or effectively optimize LiDAR placements. Moreover, these methods are limited to evaluating LiDAR placements for detection tasks since they solely rely on 3D object distribution information.
Our method addresses these limitations by introducing semantic occupancy information as a prior for the evaluation metric. This allows for more accurate characterization of boundaries in the 3D scene and effectively addresses occlusion issues between objects and the environment. Additionally, our method leverages semantic distribution information under diverse conditions, thereby enhancing the reliability of perception in adverse weather conditions and during sensor failures.
Appendix D D Additional Qualitative Result
Table 10: Comparisons of the key differences among existing sensor placement approaches.
Method Venue Deployment Sensor Prior Info Task Optimizing
S-MIG [34] CVPR 2022 Vehicle LiDAR 3D bbox Det
Jiang’s [40] ICCV 2023 Road Side LiDAR 3D bbox Det
S-MS [57] ICRA 2024 Vehicle LiDAR + Camera 3D bbox Det
Place3D Ours Vehicle LiDAR Semantic Occ Seg + Det
D.1 Clean Condition
We showcase the qualitative results of the MinkUNet [18] model using our optimized LiDAR placement. As shown in Figure 15, the depiction across the first four rows demonstrates a commendable performance of our LiDAR placement strategy during clean conditions. This is evidenced by the predominance of gray in the error maps, which indicates a high rate of correct predictions. Also, the fifth row presents a notable exception, illustrating a failure case. This particular scene exhibits a heightened complexity with a diverse array of objects and potential occlusions that challenge predictive accuracy. The qualitative results underscore the importance of continuous refinement of LiDAR placement to achieve better perception across a comprehensive spectrum of scenarios.
D.2 Adverse Conditions
In Figure 16, we showcase the qualitative results of MinkUNet [18] under different adverse conditions utilizing our optimized LiDAR placement. The corruptions from top to bottom rows are “fog”, “wet ground”, “motion blur”, “crosstalk”, and “incomplete echo”. As can be seen, the LiDAR semantic segmentation model encounters extra difficulties when predicting under such scenarios. Erroneous predictions tend to appear in regions contaminated by different types of noises, such as airborne particles, disturbed laser reflections, and jittering scatters. It becomes apparent that enhancing the model’s robustness under these adverse conditions is crucial for the practical usage of driving perception systems. To achieve better robustness, we design suitable LiDAR placements that can mitigate the degradation caused by corruptions. As discussed in Table 12, Table 13, and Table 14, our placements can largely enhance the robustness of various models.
(a) PolarNet
(b) MinkUNet
(c) SPVCNN
(d) Cylinder3D
Figure 11: The correlation between M-SOG and LiDAR semantic segmentation [18, 86, 111, 117] models performance in the clean condition.
(a) PolarNet
(b) MinkUNet
(c) SPVCNN
(d) Cylinder3D
Figure 12: The correlation between M-SOG and LiDAR semantic segmentation [18, 86, 111, 117] models performance in the adverse condition.
(a) PointPillars
(b) BEVFusion-L
(c) CenterPoint
(d) FSTR
Figure 13: The correlation between M-SOG and 3D object detection [51, 107, 63, 109] models performance in the clean condition.
(a) PointPillars
(b) BEVFusion-L
(c) CenterPoint
(d) FSTR
Figure 14: The correlation between M-SOG and 3D object detection [51, 107, 63, 109] models performance in the adverse condition.
Figure 15: Qualitative assessments of the MinkUNet [18] model using our LiDAR placement strategy. The model is tested under the clean condition. The error maps show the correct and incorrect predictions in gray and red, respectively. Kindly refer to Table 6 for color maps. Best viewed in colors and zoomed-in for details.
Figure 16: Qualitative assessments of the MinkUNet [18] model using our LiDAR placement strategy. The model is tested under adverse conditions. The error maps show the correct and incorrect predictions in gray and red, respectively. Kindly refer to Table 6 for color maps. Best viewed in colors and zoomed-in for details.
@@ D.2 Adverse Conditions
$\bullet$ Placement: Ours 2D Plane
@@
Appendix E E Proofs